# How LLMs Aid in UML Modeling: An Exploratory Study with Novice Analysts

### v1

@@ Abstract
Since the emergence of GPT-3, Large Language Models (LLMs) have caught the eyes of researchers, practitioners, and educators in the field of software engineering. However, there has been relatively little investigation regarding the performance of LLMs in assisting with requirements analysis and UML modeling. This paper explores how LLMs can assist novice analysts in creating three types of typical UML models: use case models, class diagrams, and sequence diagrams. For this purpose, we designed the modeling tasks of these three UML models for 45 undergraduate students who participated in a requirements modeling course, with the help of LLMs. By analyzing their project reports, we found that LLMs can assist undergraduate students as notice analysts in UML modeling tasks, but LLMs also have shortcomings and limitations.
@@ I Introduction
With the emergence of GPT-3, the Large Language Models (LLMs for short) have made remarkable improvements in both performance and versatility. This brings breakthroughs to natural language processing, computer vision, and various other domains. Especially, in Software Engineering (SE), LLMs have been effectively employed in code generation [1] [2], software architecture [3], software testing [4], etc. As an important and early stage in the software development life cycle, requirements engineering (RE) aims to discover, understand, formulate, analyze, and agree on what problem should be solved, why such a problem needs to be solved, and who should be involved in the responsibility of solving problems [5]. Generally, software requirements are described with natural languages or graphic models. The outstanding capability of LLMs in natural language processing and their interaction mode make them highly promising for writing requirements in natural languages. However, a few studies discussed the capability of LLMs in generating graphic requirements models. UML (Unified Modeling Language) has been widely used to help in requirements analysis and modeling with a set of graphic notations for researchers, educators, and practitioners. Therefore, it is meaningful to explore how LLMs aid in the UML modeling task for RE, and to evaluate the capability of LLMs in requirements modeling tasks.
@@
In this paper, we explored the performance of LLMs aiding in creating three types of UML models, i.e., use case diagrams, class diagrams, and sequence diagrams. More specifically, we designed and conducted an experiment in our course, i.e., software requirements analysis and modeling. 45 undergraduate students majoring in Software Engineering at Wuhan University joined in this course were asked to create UML use case diagrams, class diagrams, and sequence diagrams for a given case study [6], with the help of LLMs. By manually analyzing the created UML models as well as the human-LLM conversation for the modeling tasks in their project reports, we found that LLM can aid in the creation and optimization of these three types of UML models, but LLMs also have shortcomings and limitations. These findings may provide insights for practitioners in requirements engineering and software education professionals in the field of requirements analysis and UML modeling.
@@
The rest of this paper is organized as follows. Section II introduces the background and relevant studies on this topic. Section III defines the research questions and presents the details of our experimental design. Section IV provides the experimental results and Section V discusses these results. Section VI clarifies the limitations of this study. Section VII concludes the work and outlines the future research directions.
@@ II Related Work
Since the introduction of GPT-3, LLMs have sparked widespread interest in various academic and industrial domains. The use of LLMs in SE tasks is becoming a hot research topic [7].
@@
Many studies focused on the initial introduction and application of large language models in the field of software engineering. Ahmad et al. designed a case study that involves collaboration between a novice software architect and ChatGPT to architect a service-based software system, and the case study reflects the factors that need to be considered in collaborative architecting with ChatGPT, such as Variation in responses and artifacts, level of human decision support [3]. Luitel et al. explored whether LLMs are useful external sources of knowledge for detecting potential incompleteness in natural language requirements by utilizing BERT, and found that BERT’s predictions effectively highlight terminology that is missing from requirements [8]. White et al. provided two contributions to research on using LLMs for software engineering, a catalog of patterns for software engineering that classifies patterns according to the types of problems they solve, and several prompt patterns that have been applied to improve requirements elicitation, rapid prototyping, code quality, refactoring, and system design [9]. Zhang et al. empirically evaluated the performance of ChatGPT in requirements information retrieval (IR) tasks to derive insights into designing or developing more effective requirements retrieval methods or tools based on generative LLMs [10]. Xie et al. conducted the first empirical study to evaluate the capabilities of LLMs for generating software specifications from software comments or documentation, and their results showed that with Few Shot Learning, LLMs outperform traditional methods, and more sophisticated prompt construction strategies can further increase this performance gap [11]. Jeuring et al. explored whether computational thinking (CT) skills can predict the ability to develop software using LLM-based tools in their experiment, and the results showed that the ability to develop software using LLM-based tools can indeed be predicted by the score on a CT assessment  [12]. Waseem et al investigated how novice developers use ChatGPT during the software development life cycle (SDLC), including software design using ChatGPT and PlantUML, and the results show that ChatGPT has a positive influence on various development phases of SDLC, with enhanced efficiency, accuracy, and collaboration in software development [13]. Arora et al. explored the use of LLMs in driving RE processes, focusing on the potential for requirements elicitation, analysis, specification, and validation, and found that LLMs can enhance several RE tasks by automating, streamlining, and augmenting human capabilities [14]. Sadik et al. provided an Agile Model-Driven Development (MDD) approach to enhance code auto-generation using GPT-4, and the results show that the auto-generated code was aligned with the expected UML sequence diagram [15]. Kanuka et al. explored the feasibility of using ChatGPT to create bidirectional traceability between design models (e.g., UML class diagrams) and Code, and the study results show that ChatGPT is capable of generating design models and code from natural language requirements, and link them [16].
@@
However, in the field of requirements analysis and software modeling, especially in UML modeling, research related to LLM is still relatively scarce, and many new aspects are worth exploring and discussing. Cámara et al. investigated the current capabilities of ChatGPT to perform modeling tasks and to assist modelers, while also trying to identify its main shortcomings [17]. However, Cámara et al. only concentrated on UML class diagrams enriched with OCL constraints. Our research covers three types of UML models that are frequently used in software modeling. Furthermore, we attempt to identify the factors that influence LLM performance, such as output formats and interaction patterns, and their corresponding optimization strategies.
@@ III-B Research Questions
RQ4: Does the specified LLM output format affect the performance in UML modeling?
Rationale: In the returned project reports, some participants asked the LLM to return the specified formats, such as UML models or the PlantUML code, for the given case study. This RQ intends to explore what formats can be returned by the LLM and whether different formats of these LLM outputs affect the LLM’s performance in UML modeling. The answer to this RQ would help us understand the types of LLM output formats and which LLM output formats fit the expected UML models most.
@@ III-C1 Phase 1: Setting up modeling tasks
This phase aims to guide and conduct a specified UML modeling task aided by the LLMs, referring to Step 1 to Step 3 in Figure 1. More specifically, the participants are required to create three types of UML models, i.e., a use case model, a class diagram, and a sequence diagram, for the given case study: Order Processing System (see the case study description in [6]). Furthermore, the LLMs, such as ChatGPT, should be employed to complete this UML modeling task, to test the performance of LLM in assisting the creation of these three types of models.
The participants in this study are 45 undergraduate students majoring in Software Engineering (SE) from the School of Computer Science of our university. From February to May, 2023, they have not only learned knowledge but also got training on how to create the UML use case models, class diagrams, and sequence diagrams for a given case study during their requirements analysis and modeling course. All the participants are required to deliver their project reports in the middle of June, 2023. The report consists of two parts: (1) the UML use case model, class diagram, and sequence diagram created for the given case study in the following grey box, with the help of the selected LLM; (2) the history of the interaction between the student and the LLM.
@@
D5 Formats of UML models The notations supported by the specified tools and further used to describe the UML models in the project reports. RQ4
@@ III-C2 Phase 2: Evaluate modeling results
In this phase, we intend to analyze the collected 45 project reports, including the UML models created with the help of LLMs and the details of the human-LLM conversation. Phase 2 covers Step 4 to Step 7 in Figure  1. For this purpose, we extracted some data items from each report, as shown in Table  I.
@@
D2, D3 and D4 refer to the three types of UML models created by participants with the assistance of LLM. To evaluate the quality of these models, we defined a set of criteria to calculate the scores of the created models. Table II,  III and  IV list the evaluation criteria for the created use case model, class diagram, and sequence diagram respectively.
@@
UC1 Complete identification of actors Both user actors and system actors are included. System actors should align with the textual description, be essential in existence, and have appropriate names.
To obtain the score, core actors must be included. Different ways of expression with similar meanings are allowed.
Core actors: Customer, Gold Customer, According System
@@
To obtain the score, core use cases must be included. Different ways of expression with similar meanings are allowed.
Core use cases: Place Order, Check Order Status+Cancel Order (or Modify Order), Signup for Notification, Return Product
@@
To obtain the score, the generalization relationship between “Gold Customer” and “Customer” should be accurately represented.
UC4 Accurate recognition of relationships between actors and use cases Actors and use cases are correctly connected in a manner that aligns with the description provided in the prompt.
To obtain the score, use cases must be connected with the correct lines to actors, which is judged by modeling experts.
@@
CC1 Correct identification of classes All necessary concepts, entities, and relationships mentioned in the description are adequately covered. There are no errors or logical inconsistencies present. Class names adhere to standards and are clear and descriptive enough.
To obtain the score, core classes must be included. Different ways of expression with similar meanings are allowed.
@@
CC2 Correct identification of operations Operations are contained within the classes. The naming and design of the operations do not conflict with the class definition and generally adhere to object-oriented design principles.
To obtain the score, considering that the design concepts and structures of different class diagrams may vary significantly (while still being correct), this part will be subject to manual review by modeling experts for each class diagram.
CC3 Correct identification of attributes Attributes are contained within the classes. The naming and design of the attributes do not conflict with the class definition.
To obtain the score, considering that the design concepts and structures of different class diagrams may vary significantly (while still being correct), this part will be subject to manual review by modeling experts for each class diagram.
@@
To obtain the score, core objects must be included. Different ways of expression with similar meanings or functions are allowed.
@@
To obtain the score, even if the designs of the message system by different modelers are all reasonable, there may be significant differences. Therefore, the message part will be manually reviewed by modeling experts to ensure compliance with the formatting requirements and logic of UML modeling.
@@
To obtain the score, even if the designs of the message system by different modelers are all reasonable, there may be significant differences. Therefore, the message part will be manually reviewed by modeling experts to ensure compliance with the formatting requirements and logic of UML modeling.
@@ IV-A Demographics of the Participants
Regarding the LLMs applied in our UML modeling task, ChatGPT (including ChatGPT 3.5 and ChatGPT 4.0) is the most used one, accounting for 93.3% of the 45 participants (42 participants), as presented in Figure  3. The remaining three participants employed NewBing in their UML modeling tasks.
@@
During the human-LLM interaction, 68.9% of the 45 participants (31 participants) interacted with the LLM in English, and the remaining 14 participants used Chinese, as shown in Figure  4.
@@ IV-B Answer to RQ1: Use Case Diagram Modeling
Based on the criteria defined in Table II as well as Formula (1) and Formula (2), the UML models in all the 45 project reports were evaluated and graded. The correctness rate for each criterion regarding use case modeling and its average score is shown in Table V.
@@
As shown in Table V, all the types of modeling elements in the use case model can be identified with the help of LLMs, including actors, use cases, the relationships between actors, and the relationships between actors and user cases, etc. Particularly, the relationship between actors and use cases is the easiest to be correctly identified with the assistance of LLM, since the UML use case models in all the 45 project reports follow UC4 with a 100% correctness rate. Also, LLM performs excellently in identifying use cases, with a correctness rate of 88.89%. Surprisingly, the complete identification of the relationships between actors seems to be the most difficult task in use case modeling for the LLM, with the lowest correctness rate of 17.78%. Another challenge to LLM-aided use case modeling is to accurately identify all the actors in the given case study, as the correctness rate of UC1 is 31.11%. For example, very few participants in our study successfully identified the generalization relationship between ‘Customer’ and ‘GoldCustomer’ in the given case study.
@@ IV-D Answer to RQ3: Sequence Diagram Modeling
The correctness rate for each criterion regarding sequence diagram modeling and its average score is shown in Table VII. We found that most of the three modeling elements in the sequence diagrams can be correctly identified with the assistance of LLM. Relatively, the correct identification of the messages between objects is lower than the ones of the other two modeling elements.

### v2

@@ Abstract
Since the emergence of GPT-3, Large Language Models (LLMs) have caught the eyes of researchers, practitioners, and educators in the field of software engineering. However, there has been relatively little investigation regarding the performance of LLMs in assisting with requirements analysis and UML modeling. This paper explores how LLMs can assist novice analysts in creating three types of typical UML models: use case models, class diagrams, and sequence diagrams. For this purpose, we designed the modeling tasks of these three UML models for 45 undergraduate students who participated in a requirements modeling course, with the help of LLMs. By analyzing their project reports, we found that LLMs can assist undergraduate students as novice analysts in UML modeling tasks, but LLMs also have shortcomings and limitations that should be considered when using them.
@@ I Introduction
With the emergence of GPT-3, the Large Language Models (LLMs for short) have made remarkable improvements in both performance and versatility. This brings breakthroughs to natural language processing (NLP), computer vision, and various other domains. Especially, in Software Engineering (SE), LLMs have been effectively employed in code generation [1] [2], software architecture [3], software testing [4], etc. As an important and early stage in the software development life cycle, requirements engineering (RE) aims to discover, understand, formulate, analyze, and agree on what problem should be solved, why such a problem needs to be solved, and who should be involved in the responsibility of solving problems [5]. Generally, software requirements are described with natural languages or graphic models, and requirements modeling is a human-intensive and time-consuming task. The outstanding capability of LLMs in NLP and their interaction mode make them highly promising for writing requirements in natural languages. However, a few studies investigated the capability of LLMs in generating graphical requirements models. In SE, UML (Unified Modeling Language) has been widely used to help in requirements analysis and modeling with a set of graphic notations for researchers, educators, and practitioners. Therefore, it is meaningful to explore how LLMs aid in the UML modeling task for RE and to evaluate the capability of LLMs in requirements modeling tasks.
@@
For this purpose, this paper explored the performance of LLMs aiding in creating three types of UML models, i.e., use case diagrams, class diagrams, and sequence diagrams. More specifically, we designed and conducted an experiment in our course, i.e., software requirements analysis and modeling. 45 undergraduate students majoring in Software Engineering at Wuhan University joined in this course were asked to create UML use case diagrams, class diagrams, and sequence diagrams for a given case study [6], with the help of LLMs. By manually analyzing the created UML models as well as the human-LLM conversation for this modeling tasks in their project reports, we found that LLM can aid in the creation and optimization of these three types of UML models, but LLMs still have some shortcomings and limitations. These findings may provide insights for practitioners in requirements engineering and software education professionals in the field of requirements analysis and UML modeling.
The main contributions of this paper are as follows.
  • The effectiveness of LLMs in creating three types of UML models, i.e. use case diagrams, class diagrams, and sequence diagrams, was investigated and evaluated with the course projects in the undergraduate program in university.
  • The evaluation criteria for the created three types of UML models with LLMs were proposed in detail to cover the core modeling elements of these three UML models.
  • The output formats of the LLM-aided UML models were summarized to explore the impact factors of applying LLMs in the UML modeling tasks.
@@
The remainder of this paper is organized as follows. Section II introduces the background and relevant studies on this topic. Section III defines the research questions and presents the details of our experimental design. Section IV provides the experimental results, and Section V discusses these results. Section VI addresses the limitations of this study. Section VII concludes the work and future research directions.
@@ II Related Work
With the development of GPT-3, LLMs have sparked widespread interest in various academic and industrial domains. The use of LLMs in SE tasks is now becoming a hot research topic [7].
@@
In the SE community, researchers have conducted some studies using large language models, primarily focusing on code generation and software architecture assistance. For example, Ahmad et al. designed a case study that involves collaboration between a novice software architect and ChatGPT to architect a service-based software system. The case study reflects the factors that need to be considered in collaborative architecture design with ChatGPT, such as Variation in responses and artifacts, level of human decision support [3]. White et al. provided a catalog of patterns for software engineering that classifies patterns according to the types of problems they solve. They also explored several prompt patterns that had been applied to improve requirements elicitation, rapid prototyping, code quality, refactoring, and system design [8]. Xie et al. conducted the first empirical study to evaluate the capabilities of LLMs to generate software specifications from software comments or documentation. The results showed that with few shot learning, LLMs outperformed the traditional methods and more sophisticated prompt construction strategies can further increase this performance gap [9]. Jeuring et al. explored whether computational thinking (CT) skills can predict the ability to develop software using LLM-based tools in their experiment, and the results showed that the ability to develop software using LLM-based tools can indeed be predicted by the score on a CT assessment  [10]. Sadik et al. provided an Agile Model-Driven Development (MDD) approach to enhance code auto-generation using GPT-4, and the results show that the auto-generated code was aligned with the expected UML sequence diagram [11].
As for RE, there is also an amount of researches about the application of LLMs across various stages of the RE process. Luitel et al. explored whether LLMs are useful external sources of knowledge for the detection of potential incompleteness in natural language requirements by utilizing BERT. They found that BERT-based predictions effectively highlight terminology that is missing from requirements descriptions [12]. Zhang et al. empirically evaluated the performance of ChatGPT in requirements information retrieval (IR) tasks to derive insights into the design or development of more effective requirements retrieval methods or tools based on generative LLMs [13]. Waseem et al. investigated how novice developers use ChatGPT during the software development life cycle (SDLC), including software design using ChatGPT and PlantUML. The results revealed that ChatGPT has a positive influence on various development phases of SDLC, with enhanced efficiency, accuracy, and collaboration in software development [14]. Arora et al. explored the use of LLMs in driving RE processes, focusing on the potential for requirements elicitation, analysis, specification, and validation, and found that LLMs can enhance several RE tasks by automating, streamlining, and augmenting human capabilities [15]. Kanuka et al. explored the feasibility of using ChatGPT to create bidirectional traceability between design models (e.g., UML class diagrams) and Code, and the study results show that ChatGPT is capable of generating design models and code from natural language requirements, and link them [16].
@@
However, only a few researcher employed LLMs in requirements analysis and modeling, especially in UML modeling. For instance, Cámara et al. investigated the strengthens and weakness of ChatGPT in helping with modeling tasks [17]. However, the authors only concentrated on UML class diagrams enriched with OCL constraints. Our work covers three types of UML models that are frequently used in requirements modeling. Furthermore, we attempt to identify the factors that influence LLM performance in the requirements modeling tasks with UML, such as output formats.
@@ III-B Research Questions
RQ4: Does the specified output format affect the performance of LLMs in UML modeling?
Rationale: In the project reports, some participants asked the LLM to return the specified formats, such as UML models or the PlantUML code, for the given case study. This RQ intends to explore what formats can be returned by the LLM and whether different formats of these LLM outputs affect the LLM’s performance in UML modeling. The answer to this RQ would help us understand the types of LLM output formats and which LLM output formats fit the expected UML models most.
@@ III-C1 Phase 1: Setting up modeling tasks
This phase aims to guide and conduct a specified UML modeling task aided by LLMs, referring to Step 1 to Step 3 in Figure 1. More specifically, the participants are required to create three types of UML models, i.e., a use case model, a class diagram, and a sequence diagram, for the given case study: Order Processing System (see the case study description in [6]). Furthermore, the LLMs, such as ChatGPT, should be employed to complete this UML modeling task, in order to examine the performance of LLM in assisting the creation of these three types of models.
The participants in this study are 45 undergraduate students majoring in Software Engineering (SE) from the School of Computer Science of our university. From February to May, 2023, they have not only learned knowledge but also got training on how to create the UML use case models, class diagrams, and sequence diagrams for a given case study during their requirements analysis and modeling course. All the participants are required to deliver their project reports in the middle of June, 2023. The report consists of two parts: (1) the UML use case model, class diagram, and sequence diagram created for the given case study in [6], with the help of the selected LLM; (2) the history of the interaction between the student and the LLM.
@@
D5 Formats of UML models The notations supported by the specified tools and used to describe the UML models in the project reports. RQ4
@@ III-C2 Phase 2: Evaluate modeling results
In this phase, we intend to analyze the collected 45 project reports, including the UML models created with the help of LLMs and the details of the human-LLM conversation. Phase 2 covers Step 4 to Step 7 in Figure  1. To evaluate the modeling results, we extracted some data items from each report, as shown in Table  I.
@@
D2, D3, and D4 refer to the three types of UML models created by participants with the assistance of LLM. To evaluate the quality of these models, we defined a set of criteria to calculate the scores of the created models. Table II,  III and  IV list the evaluation criteria for the created use case model, class diagram, and sequence diagram respectively.
@@
UC1 Complete identification of actors Primary actors should align with the textual description, be essential in existence, and have appropriate names.
To obtain the score, primary actors must be included. Different expressions with similar meanings are allowed.
Primary actors: Customer, Gold Customer, Accounting System
@@
To obtain the score, core use cases must be included. Different expressions with similar meanings are allowed.
Core use cases: Place Order, Check Order Status, Cancel Order (or Modify Order), Signup for Notification, Return Product
@@
To obtain the score, the generalization relationship between ‘Gold Customer’ and ‘Customer’ should be accurately represented.
UC4 Accurate recognition of relationships between actors and use cases Actors and use cases are correctly connected, aligning with the description provided in the prompt.
To obtain the score, use cases must be connected to the corresponding actors, with the correct UML notations.
@@
CC1 Correct identification of classes All necessary concepts and relationships mentioned in the description are adequately covered. There are no errors or logical inconsistencies. The names of classes are clear and descriptive enough.
To obtain the score, core classes must be included. Different expressions with similar meanings are allowed.
@@
CC2 Correct identification of operations Operations are contained in the corresponding classes. The included operations should be named appropriately to avoid conflicts with the class and generally adhere to object-oriented design principles.
To obtain the score, the class diagram should well-structured and reasonable. It will be manually reviewed and evaluated by the authors of this paper.
CC3 Correct identification of attributes Attributes are contained within the classes. The names of the attributes do not conflict with the class.
To obtain the score, the attributes should be defined with appropriate names and fit the class they belong to. It will be manually reviewed and evaluated by the authors of this paper.
@@
To obtain the score, core objects must be included. Different expressions with similar meanings or functions are allowed.
@@
To obtain the score, the source and target objects of the message as well as the names of the messages should be reasonable. The identification of messages will be manually reviewed and evaluated by the authors of this paper.
@@
To obtain the score, the order of the messages should be reasonable. The order of messages will be manually reviewed and evaluated by the authors of this paper.
@@
Figure 2: Examples of the three output formats of LLMs
@@ IV-A Demographics of the Participants
Regarding the LLMs applied in our UML modeling task, ChatGPT (including ChatGPT 3.5 and ChatGPT 4.0) is the most used one, accounting for 93.3% of the 45 participants (42 participants), as presented in Figure  4. The remaining three participants employed NewBing in their UML modeling tasks.
@@
During the human-LLM interaction, 68.9% of the 45 participants (31 participants) interacted with the LLM in English, and the remaining 14 participants used Chinese, as shown in Figure  5.
@@ IV-B Answer to RQ1: Use Case Diagram Modeling
Based on the criteria defined in Table II as well as Formula (1) and Formula (2), the UML models in the 45 project reports were evaluated and graded. The correctness rate for each criterion regarding use case modeling and its average score is shown in Table V.
@@
As shown in Table V, all types of modeling elements in the use case model can be identified with the help of LLM, including actors, use cases, the relationships between actors, and the relationships between actors and user cases, etc. Particularly, the relationship between actors and use cases is the easiest to be correctly identified with the assistance of LLM, since the UML use case models in all the 45 project reports follow UC4 with a 100% correctness rate. In addition, LLM performs excellently in identifying use cases, with a correctness rate of 88.89%. Surprisingly, the complete identification of the relationships between actors seems to be the most difficult task in use case modeling for the LLM, with the lowest correctness rate of 17.78%. Another challenge to LLM-aided use case modeling is to accurately identify all the actors in the given case study, as the correctness rate of UC1 is 31.11%. For example, very few participants in our study successfully identified the generalization relationship between ‘Customer’ and ‘GoldCustomer’ in the given case study.
@@ IV-D Answer to RQ3: Sequence Diagram Modeling
The correctness rate for each criterion with respect to sequence diagram modeling and its average score is shown in Table VII. We found that most of the three modeling elements in the sequence diagrams can be correctly identified with the help of LLM. Relatively, the correct identification of the messages between objects is lower than those of the other two modeling elements.